OpenStudy (anonymous):

which statements are true? choose all that are correct? A. 7+(-10)<-13 A. 7 + (–10) < –13 + 17 B. 22 + (–13) > 18 + (–7) C. –18 + 24 = –23 + 17 D. –11 + (–8) > 4 + (–26)

jimthompson5910 (jim_thompson5910):

what is 7+(-10) equal to

OpenStudy (anonymous):

I don't know thats why I asked someone for help

jimthompson5910 (jim_thompson5910):

draw out a number line then plot 7 on this number line afterwards, go to the left 10 units. Where do you end up?

jimthompson5910 (jim_thompson5910):

here's a number line http://www.watertown.k12.ma.us/wms/math/math_help/images/atnnumberline.JPEG

OpenStudy (anonymous):

um -3????

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so 7+(-10) = -3

jimthompson5910 (jim_thompson5910):

you can calculate that without a number line by saying 10 - 7 = 3 the answer will be negative because the 10 is bigger than 7 and the 10 is negative

jimthompson5910 (jim_thompson5910):

what is -13 + 17 equal to?

OpenStudy (anonymous):

um im only in fifth grade and im sick so I missed the teacher explaining how to do it so if you would please break it down a little more for me please

jimthompson5910 (jim_thompson5910):

which step?

jimthompson5910 (jim_thompson5910):

you see how 7+(-10) = -3 right?

OpenStudy (anonymous):

the what is -3+17 is equal to what

jimthompson5910 (jim_thompson5910):

-13+17 is the same as 17 + (-13) I flipped the positions of -13 and 17. I'm still adding.

jimthompson5910 (jim_thompson5910):

17 + (-13) is the same as 17 - 13

OpenStudy (anonymous):

so the answer is 4

jimthompson5910 (jim_thompson5910):

yes, -13 + 17 = 4

jimthompson5910 (jim_thompson5910):

so 7 + (-10) < -13 + 17 turns into -3 < 4 is that second inequality true?

OpenStudy (anonymous):

yes???

jimthompson5910 (jim_thompson5910):

yes because -3 is definitely less than 4 on the number line, -3 is to the left of 4, so -3 < 4 is true

jimthompson5910 (jim_thompson5910):

because -3 < 4 is true this means 7 + (-10) < -13 + 17 is also true

jimthompson5910 (jim_thompson5910):

you will do the same for parts B through D

OpenStudy (anonymous):

oh ok thank you and I have a few more questions can you help me answer them???

jimthompson5910 (jim_thompson5910):

I can do 2 more

OpenStudy (anonymous):

ok thank you and what grade are you in

jimthompson5910 (jim_thompson5910):

I'm in college